### v2

# Random Projection Layers for Multidimensional Time Series Forecasting

Abstract
All-Multi-Layer Perceptron (all-MLP) mixer models have been shown to be effective for time series forecasting problems. However, when such a model is applied to high-dimensional time series (e.g., the time series in a spatial-temporal dataset), its performance is likely to degrade due to overfitting issues. In this paper, we propose an all-MLP time series forecasting architecture, referred to as RPMixer. Our method leverages the ensemble-like behavior of deep neural networks, where each individual block within the network acts like a base learner in an ensemble model, especially when identity mapping residual connections are incorporated. By integrating random projection layers into our model, we increase the diversity among the blocks’ outputs, thereby enhancing the overall performance of RPMixer. Extensive experiments conducted on large-scale spatial-temporal forecasting benchmark datasets demonstrate that our proposed method outperforms alternative methods, including both spatial-temporal graph models and general forecasting models.
1 Introduction
Multidimensional time series forecasting systems play a crucial role in addressing various real-world problems Yan et al. (2018); Yao et al. (2018); Geng et al. (2019); Liu et al. (2020); Cai et al. (2023). Over the years, a multitude of multidimensional time series forecasting methods have been proposed Shang et al. (2021); Zeng et al. (2023); Chen et al. (2023). In this paper, we have a particular interest in solutions that exclusively utilize Multi-Layer Perceptron (MLP) due to its simplicity, efficiency, and state-of-the-art performance in time series forecasting Chen et al. (2023). These models, often dubbed all-MLP or mixer models, consist of layers of mixer blocks Tolstikhin et al. (2021); Chen et al. (2023). However, when applying existing methods like TSMixer Chen et al. (2023) to high-dimensional datasets, such as the multidimensional time series from spatial-temporal datasets, the performance tends to suffer, likely due to overfitting.
Our proposed method seeks to enhance the existing mixer model’s capacity to capture the relationship between different dimensions of the input time series. Specifically, we incorporate random projection into mixer models to reinforce the model’s ability to learn such relationships. It has been observed that deep learning models exhibit ensemble-like behavior when identity mapping connections are integrated into the model design Veit et al. (2016). This results in residual units, or mixer blocks in the context of mixer models, functioning similarly to base learners in an ensemble model Veit et al. (2016). Given that diversity is a critical factor for the success of an ensemble model Margineantu and Dietterich (1997), and diversifying the intermediate outputs is akin to diversifying the base learners in an ensemble model, our proposed method has demonstrated superior performance compared to existing solutions on large-scale spatial-temporal forecasting benchmark datasets.
Figure 1: Each row consists of the intermediate and final outputs from a sampled node, while each column consists of the outputs from different mixer blocks within the model. The three samples are from node 238, 280, and 295 of the SD dataset (see Section 5 for details).
To illustrate that the inclusion of random projection does diversify the intermediate representation, we refer to Fig. 1, which presents three examples of the outputs from the first, third, fifth, seventh, and final (eighth) blocks of two mixer models. One of these models integrates random projection (i.e., “proposed" in Fig. 1), while the other does not. As shown, the outputs of the mixer blocks in our proposed method exhibit greater diversity (i.e., the outputs from different blocks are less similar) compared to those of alternative models that lack random projection. This diversification leads to more accurate predictions, as evidenced by the output of the final block, which closely mirrors the ground truth. Due to our application of random projection in the mixer model design, we have aptly named our proposed method RPMixer.
The key contributions of this paper include:
  • We have developed a novel spatial-temporal forecasting method, RPMixer, which is both effective and efficient. We incorporate random projection layers into RPMixer to enhance the modeling diversity and capability of mixer models.
  • Our ablation study highlights the importance of identity mapping for achieving the best performance of our proposed RPMixer.
  • Extensive experiments validate the effectiveness of our proposed method on large-scale spatial-temporal forecasting datasets, demonstrating superior performance compared to existing solutions.
2 Related Work
In this section, we review the literature on three areas: 1) spatial-temporal forecasting methods, 2) multidimensional time series forecasting methods, and 3) random projection methods.
[…]
The spatial-temporal forecasting problem, in the absence of graph input, essentially transforms into a multidimensional time series forecasting problem (see Section 3 for details). As such, we also review methods designed for multidimensional time series forecasting. Over the years, transformer-based methods Vaswani et al. (2017) like LogTrans Li et al. (2019), Pyraformer Liu et al. (2021), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and Fedformer Zhou et al. (2022) have emerged. However, simple linear models outperform these transformer-based methods on multidimensional long-term time series forecasting datasets as Zeng et al. (2023) demonstrates. In contrast, mixer-based methods Tolstikhin et al. (2021), like TSMixer Chen et al. (2023), have demonstrated promising performance on multidimensional time series forecasting problems. Consequently, we aim to develop a mixer-based architecture for the spatial-temporal forecasting problem.
Random projection Bingham and Mannila (2001) is utilized in a variety of machine learning and data mining methods Bingham and Mannila (2001); Schclar and Rokach (2009); Cannings and Samworth (2017); Yeh et al. (2022a). However, the majority of these works Bingham and Mannila (2001); Cannings and Samworth (2017); Yeh et al. (2022a) concentrate on the efficiency gains from random projection. Only Schclar and Rokach (2009) discusses how the diversity introduced by random projection could enhance the performance of an ensemble model, but without confirming that random projection actually introduces diversity into the model. Besides random projection, randomized methods like random shapelets/convolutions Renard et al. (2016); Yeh and Keogh (2016); Yeh (2018); Dempster et al. (2020) also perform exceptionally well for time series data on classification problems. Some papers also employ neural networks with partially fixed random initialized weights Huang et al. (2015); Ulyanov et al. (2018); Rosenfeld and Tsotsos (2019), but these papers neither focus on the spatial-temporal forecasting problem nor provide an analysis based on the ensemble interpretation of deep neural networks Veit et al. (2016). To the best of our knowledge, our paper is the first to investigate fixed randomized layers (or random projection) in the context of spatial-temporal forecasting problems.
3 Definition and Problem
We use lowercase letters (e.g., $x$), boldface lowercase letters (e.g., $\mathbf{x}$), uppercase letters (e.g., $X$), and boldface uppercase letters (e.g., $\mathbf{X}$) to denote scalars, vectors, matrices, and tensors, respectively. We begin by introducing the components of a spatial-temporal dataset. The spatial and temporal information are stored in the adjacency matrix and the time series matrix, respectively.
Definition 1.
Given that there are $n$ entities in a spatial-temporal dataset, an adjacency matrix $A\in\mathbb{R}^{n\times n}$ stores the spatial relationships among the entities, i.e., $A[i,j]$ describes the relationship between the $i$-th entity and the $j$-th entity.
Definition 2.
Given that there are $n$ entities in a spatial-temporal dataset, a time series matrix $X\in\mathbb{R}^{n\times t}$ stores the temporal information of the entities, where $t$ is the length of the time series in the dataset.
Next, we introduce the spatial-temporal forecasting problem.
Problem 1.
Given an adjacency matrix $A\in\mathbb{R}^{n\times n}$ and a historical time series matrix $X_{\text{past}}\in\mathbb{R}^{n\times t_{\text{past}}}$ for the past $t_{\text{past}}$ steps, the goal of spatial-temporal forecasting is to learn a model $F(\cdot)$ which predicts the future time series matrix $X_{\text{future}}\in\mathbb{R}^{n\times t_{\text{future}}}$ for the next $t_{\text{future}}$ steps. The problem can be formulated as: $F(X_{\text{past}},A)\rightarrow X_{\text{future}}$.
The problem formulation aligns with the multidimensional time series forecasting problem [^1] Each node in the spatial-temporal data is considered a dimension of a multidimensional time series. Zeng et al. (2023) if the adjacency matrix $A$ is disregarded by the model $F(\cdot)$. It is important to note that each node in a spatial-temporal dataset can be associated with multiple feature dimensions. This means that the input time series to the model $F(\cdot)$ could become a time series tensor $\mathbf{X}_{\text{past}}\in\mathbb{R}^{n\times d\times t_{\text{past}}}$, where $d$ represents the number of features of the time series and $t_{\text{past}}$ is the number of time steps in the input time series. Linear or MLP-based models can easily accommodate this additional dimensionality by reshaping the time series tensor $\mathbf{X}_{\text{past}}$ into a matrix with the shape $n\times dt_{\text{past}}$ as they utilize fully connected (linear) layers. Given the $dt_{\text{past}}$-sized vector associated with an entity, a fully connected layer is capable of utilizing information from every dimension in each past time step for the prediction.
4 Proposed Method
The proposed RPMixer adopts an all-MLP mixer architecture akin to the TSMixer Chen et al. (2023) method, but it differs in three aspects:
  1. 1.
    Our design emphasizes the inclusion of identity mapping connections, achieved with a pre-activation design. The RPMixer is inspired by the ensemble interpretation of the residual connection Veit et al. (2016), which necessitates the incorporation of identity mapping connections in the design.
  2. 2.
    We incorporate random projection layers in our model to increase diversity among the outputs of different mixer blocks. This is based on the ensemble interpretation, where the mixer blocks serve a similar function to the base learners in an ensemble model.
  3. 3.
    We process the time series in the frequency domain using complex linear layers, as time series generated from human activity are typically periodic. However, for non-periodic time series, complex layers may not be optimal. The determination of how periodic a signal should be to justify the use of complex linear layers is a topic for future research.
The overall model design is depicted in Fig. 2. Given an input historical time series matrix $X_{\text{past}}\in\mathbb{R}^{n\times t_{\text{past}}}$, the model processes it with $n_{\text{block}}$ mixer blocks, which we will introduce in the next paragraph. The output of each mixer block is also in $\mathbb{R}^{n\times t_{\text{past}}}$. The output of the last mixer block is processed by an output linear layer to project the length of the time series to the desired length (i.e., $t_{\text{future}}$). The output time series matrix is in $\mathbb{R}^{n\times t_{\text{future}}}$. We optimize the model with mean absolute error (MAE) loss.
Figure 2: The RPMixer architecture.
The design of the mixer block is illustrated in Fig. 3. The highlighted forward path (in red) is used to explain the identity mapping connection in Section 4.2. The shape of the input, intermediate, and output matrices are also included in the figure. The sizes of the input and output matrices are both in $\mathbb{R}^{n\times t_{\text{past}}}$.
Figure 3: The detailed design of the mixer block. The red line highlighted the identity mapping connection for a mixer block.
The mixer block comprises two sub-blocks: 1) a temporal mixer block focusing on modeling the relationship between different time steps, and 2) a spatial mixer block focusing on modeling the relationship between different nodes. The temporal mixer block employs a complex linear layer after the ReLU activation function to model the time series in the frequency domain. We delve into the details of the complex linear layer in Appendix A.1.
The first and last operations of the spatial mixer block are matrix transpositions. This is done because the linear layers within the spatial mixer block are designed to model the relationship between the nodes dimension of the matrix. By transposing the matrix, we enable the linear layers to linearly combine the representations associated with each node, as opposed to time steps.
There are two types of linear layers in the spatial mixer block: 1) the random projection layer, and 2) the regular linear layer. We first use the random projection layer to project the vectors within the matrix from $\mathbb{R}^{n}$ to $\mathbb{R}^{n_{\text{rand}}}$. We regard the random projection layer as a type of linear layer because it consists of linear operations, which we implemented with a fixed, randomly initialized linear layer. We consider $n_{\text{rand}}$ a hyper-parameter of our method and provide an empirical study of this parameter in Section 5.5. The details about the random projection layer are discussed in Section 4.1. Subsequently, a regular linear layer is used to map the size of the vector within the matrix from $n_{\text{rand}}$ back to $n$. Both linear layers are preceded by a ReLU activation function.
4.1 Random Projection Layer
With the ensemble interpretation established in Section 4.2, our goal is to utilize this interpretation to further enhance the performance of our mixer model. Specifically, we aim to increase the diversity among the outputs of different base learners by incorporating random projection layers.
{python}
import torch import torch.nn as nn import torch.nn.functional as F
class RPLayer(nn.Module): def __init__(self, in_dim, out_dim, seed): super(RPLayer, self).__init__() torch.manual_seed(seed=seed) weight = torch.randn( out_dim, in_dim, requires_grad=False) self.register_buffer( ’weight’, weight, persistent=True) self.register_buffer( ’bias’, None)
def forward(self, x): return F.linear(x, self.weight, self.bias)
As demonstrated in the above pseudocode, the random projection layer is a fixed, randomly initialized linear layer that performs random projection Bingham and Mannila (2001) on the input. If we split the input time series matrix $X\in\mathbb{R}^{n\times t_{\text{past}}}$ into $t_{\text{past}}$ vectors of size $n$, the random projection layer computes $n_{\text{rand}}$ random combinations of the vector elements to form new vectors of size $n_{\text{rand}}$. In essence, by keeping the weights of the random projection layers fixed during training, we encourage each mixer block to concentrate on different nodes, since each random projection layer possesses its own set of randomly initialized weights.
Figure 4: The time series output from a random projection layer retains the daily and weekly patterns of the original time series.
According to the Johnson-Lindenstrauss lemma Johnson (1984), the relationships between the $t_{\text{past}}$ vectors are preserved after the random projection layer with high probability. This implies that the dynamics of the overall spatial-temporal data are largely retained in the data, even if it is projected to a much smaller space. To visually verify this claim, we randomly sampled nine nodes from the LargeST dataset Liu et al. (2023) to form a nine-dimensional time series and used random projection to reduce the dimensionality to three. The visualization is shown in Fig. 4, where the daily and weekly patterns are preserved in the time series output by random projection. A similar phenomenon has been observed in Yeh et al. (2023b) for the time series discord mining problem.
Another advantage of the random projection layer is its efficiency. As a dimension reduction technique Bingham and Mannila (2001), random projection can improve the model’s efficiency, especially when dealing with spatial-temporal data with a large number of nodes, by reducing the vector size from $n$ to a smaller $n_{\text{rand}}$. Overall, the random projection layers promote the diversity among each mixer block’s outputs while ensure efficiency.
4.2 Identity Mapping with Pre-Activation
The identity mapping connection He et al. (2016) plays an important role in the ensemble interpretation Veit et al. (2016) of the proposed RPMixer. It facilitates the creation of shorter paths, which acts like base learners within the model. We have implemented the identity mapping connection drawing inspiration from the pre-activation design proposed in He et al. (2016), where the activation functions (e.g., ReLU) precede the weighted layers.
To understand why shorter paths exist within the model, we refer back to Fig. 3. We define the weighted paths for the temporal and spatial mixer sub-blocks as follows:
$F_{\text{temp}}(X)\vcentcolon=\textsc{ComplexLinear}(\textsc{ReLU}(X))$ (1)
$F_{\text{sp}}(X)\vcentcolon=\textsc{Linear}(\textsc{ReLU}(\textsc{RandProject}(\textsc{ReLU}(X^{T}))))^{T}$ (2)
Given an input $X$, the operation of the mixer block can be expressed as:
$\textsc{Mixer}(X)=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)+X$ (3)
If we further define a function $G(\cdot)$ as:
$G(X)\vcentcolon=F_{\text{sp}}(F_{\text{temp}}(X)+X)+F_{\text{temp}}(X)$ (4)
The mixer block operation can be simplified to:
$\textsc{Mixer}(X)=G(X)+X$ (5)
Figure 5: An example used for illustrating how the RPMixer could be interpreted as an ensemble model. In this example, there are only three mixer blocks in the model; however, the same analysis can be extended to models with more blocks.
In this equation, the first term represents the weighted path and the second term is the identity mapping connection (as shown by the red line in Fig. 3). With this simplified notation in place, let us consider the case depicted in Fig. 5, which illustrates the RPMixer with three mixer blocks. Here, we use $D(\cdot)$ to represent the output linear layer. Note, the example presented in the figure only consists of three mixer blocks, but the same analysis can be extended to models with more mixer blocks.
Figure 6: The unraveled view of different forward passes for the model illustrated in Fig. 5.
Following the analysis presented in Veit et al. (2016), we can unravel the forward pass in Fig. 5 into multiple paths. The unraveled view is illustrated in Fig. 6.
If we denote the outputs of $G_{1}(\cdot)$, $G_{2}(\cdot)$, and $G_{3}(\cdot)$ as $H_{1}$, $H_{2}$, and $H_{3}$ respectively, the output $Y$ can be represented as: $Y=D(X+H_{1}+H_{2}+H_{3})$. Since $D(\cdot)$ is a linear function, the above equation can be rewritten as: $Y=D(X)+D(H_{1})+D(H_{2})+D(H_{3})$. If we further define: $Y_{0}\vcentcolon=D(X)$, $Y_{1}\vcentcolon=D(H_{1})$, $Y_{2}\vcentcolon=D(H_{2})$, and $Y_{3}\vcentcolon=D(H_{3})$, it becomes clear that the prediction $Y$ is the sum of the individual predictions from each path, i.e., $Y=Y_{0}+Y_{1}+Y_{2}+Y_{3}$. In essence, the identity mapping connections introduced by the pre-activation design facilitate an ensemble-like behavior in RPMixer.
5 Experiment
In this section, we present experiment results that demonstrate the effectiveness of our proposed method. We begin by introducing the dataset, benchmark settings, and baseline methods. Following this, we explore the benchmark results and an ablation study, which showcases the impact of each design choice. Importantly, we show how the random projection layer significantly enhances the diversity of the network’s intermediate representation, contributing substantially to the superior performance of our proposed method. Next, we conduct a sensitivity analysis on the crucial hyper-parameters. It is worth noting that our proposed method is a general approach for multivariate time series forecasting. Thus, we also perform experiments on benchmark datasets for multivariate time series forecasting and include the results in Appendix A.6. Details regarding the implementation can be found in Appendix A.4. The source code for our experiments can be downloaded from The Author(2023) (s).
5.1 Dataset and Benchmark Setting
We conducted our experiments using the LargeST dataset Liu et al. (2023), which consists of traffic data collected from 8,600 sensors in California from 2017 to 2021. For benchmarking purposes, we generated four sub-datasets following the procedure outlined in Liu et al. (2023). These sub-datasets, namely SD, GBA, GLA, and CA, include sensor data from the San Diego region, the Greater Bay Area region, the Greater Los Angeles region, and all 8,600 sensors across California, respectively. The statistics pertaining to these datasets are detailed in Appendix A.2.
In line with the experimental setup described in Liu et al. (2023), we only utilized traffic data from 2019. Sensor readings, originally recorded at 5-minute intervals, were aggregated into 15-minute windows, yielding 96 windows per day. Each sub-dataset was chronologically divided into training, validation, and test sets at a ratio of 6:2:2. The benchmark task was to predict the next 12 steps for each sensor at each timestamp. Performance was measured using the mean absolute error (MAE), root mean squared error (RMSE), and mean absolute percentage error (MAPE).
5.2 Baseline Method
[…]
Table 1: Performance comparisons. We bold the best-performing results. The performance reported in the “Average" column is computed by averaging over 12 predicted time steps. The absence of baselines on the GLA and CA datasets indicates that the models incur an out-of-memory issue. Param: the number of learnable parameters. K: $10^{3}$. M: $10^{6}$.
Data Method Param Horizon 3 Horizon 6 Horizon 12 Average
MAE RMSE MAPE MAE RMSE MAPE MAE RMSE MAPE MAE RMSE MAPE
SD HL 33.61 50.97 20.77% 57.80 84.92 37.73% 101.74 140.14 76.84% 60.79 87.40 41.88%
LSTM 98K 19.17 30.75 11.85% 26.11 41.28 16.53% 38.06 59.63 25.07% 26.73 42.14 17.17%
ASTGCN 2.2M 20.09 32.13 13.61% 25.58 40.41 17.44% 32.86 52.05 26.00% 25.10 39.91 18.05%
DCRNN 373K 17.01 27.33 10.96% 20.80 33.03 13.72% 26.77 42.49 18.57% 20.86 33.13 13.94%
AGCRN 761K 16.05 28.78 11.74% 18.37 32.44 13.37% 22.12 40.37 16.63% 18.43 32.97 13.51%
STGCN 508K 18.23 30.60 13.75% 20.34 34.42 15.10% 23.56 41.70 17.08% 20.35 34.70 15.13%
GWNET 311K 15.49 25.45 9.90% 18.17 30.16 11.98% 22.18 37.82 15.41% 18.12 30.21 12.08%
STGODE 729K 16.76 27.26 10.95% 19.79 32.91 13.18% 23.60 41.32 16.60% 19.52 32.76 13.22%
DSTAGNN 3.9M 17.83 28.60 11.08% 21.95 35.37 14.55% 26.83 46.39 19.62% 21.52 35.67 14.52%
DGCRN 243K 15.24 25.46 10.09% 17.66 29.65 11.77% 21.38 36.67 14.75% 17.65 29.70 11.89%
[…]
1NN - 21.79 35.15 13.79% 25.64 41.59 17.05% 30.77 50.59 22.38% 25.47 41.36 17.25%
Linear 3.5K 20.58 33.30 12.98% 27.00 43.87 18.20% 32.35 53.51 22.38% 25.85 42.35 17.10%
TSMixer 815K 17.13 27.42 11.35% 19.30 31.07 12.50% 22.03 35.70 14.26% 19.06 30.66 12.55%
RPMixer 1.5M 15.12 24.83 9.97% 17.04 28.24 10.98% 19.60 32.96 13.12% 16.90 27.97 11.07%
GBA HL 32.57 48.42 22.78% 53.79 77.08 43.01% 92.64 126.22 92.85% 56.44 79.82 48.87%
LSTM 98K 20.41 33.47 15.60% 27.50 43.64 23.25% 38.85 60.46 37.47% 27.88 44.23 24.31%
ASTGCN 22.3M 21.40 33.61 17.65% 26.70 40.75 24.02% 33.64 51.21 31.15% 26.15 40.25 23.29%
DCRNN 373K 18.25 29.73 14.37% 22.25 35.04 19.82% 28.68 44.39 28.69% 22.35 35.26 20.15%
AGCRN  [^7] K 18.11 30.19 13.64% 20.86 34.42 16.24% 24.06 39.47 19.29% 20.55 33.91 16.06%
STGCN 1.3M 20.62 33.81 15.84% 23.19 37.96 18.09% 26.53 43.88 21.77% 23.03 37.82 18.20%
GWNET 344K 17.74 28.92 14.37% 20.98 33.50 17.77% 25.39 40.30 22.99% 20.78 33.32 17.76%
STGODE 788K 18.80 30.53 15.67% 22.19 35.91 18.54% 26.27 43.07 22.71% 21.86 35.57 18.33%
DSTAGNN 26.9M 19.87 31.54 16.85% 23.89 38.11 19.53% 28.48 44.65 24.65% 23.39 37.07 19.58%
DGCRN 374K 18.09 29.27 15.32% 21.18 33.78 18.59% 25.73 40.88 23.67% 21.10 33.76 18.58%
[…]
1NN - 24.84 41.30 17.70% 29.31 48.56 22.92% 35.22 58.44 31.07% 29.10 48.23 23.14%
Linear 3.5K 21.55 34.79 17.94% 27.24 43.36 23.66% 31.50 51.56 26.18% 26.12 42.14 22.10%
TSMixer 3.1M 17.57 29.22 14.14% 19.85 32.64 16.95% 22.27 37.60 18.63% 19.58 32.56 16.58%
RPMixer 2.3M 17.35 28.69 13.42% 19.44 32.04 15.61% 21.65 36.20 17.42% 19.06 31.54 15.09%
GLA HL 33.66 50.91 19.16% 56.88 83.54 34.85% 98.45 137.52 71.14% 59.58 86.19 38.76%
LSTM 98K 20.09 32.41 11.82% 27.80 44.10 16.52% 39.61 61.57 25.63% 28.12 44.40 17.31%
ASTGCN 59.1M 21.11 34.04 12.29% 28.65 44.67 17.79% 39.39 59.31 28.03% 28.44 44.13 18.62%
DCRNN 373K 18.33 29.13 10.78% 22.70 35.55 13.74% 29.45 45.88 18.87% 22.73 35.65 13.97%
AGCRN 792K 17.57 30.83 10.86% 20.79 36.09 13.11% 25.01 44.82 16.11% 20.61 36.23 12.99%
STGCN 2.1M 19.87 34.01 12.58% 22.54 38.57 13.94% 26.48 45.61 16.92% 22.48 38.55 14.15%
GWNET 374K 17.30 27.72 10.69% 21.22 33.64 13.48% 27.25 43.03 18.49% 21.23 33.68 13.72%
STGODE 841K 18.46 30.05 11.94% 22.24 36.68 14.67% 27.14 45.38 19.12% 22.02 36.34 14.93%
DSTAGNN 66.3M 19.35 30.55 11.33% 24.22 38.19 15.90% 30.32 48.37 23.51% 23.87 37.88 15.36%
1NN - 23.23 38.69 13.44% 27.75 45.92 17.07% 33.49 55.51 22.86% 27.49 45.57 17.28%
Linear 3.5K 21.32 34.48 13.35% 27.45 43.83 17.79% 32.50 52.69 21.76% 26.40 42.56 17.16%
TSMixer 4.6M 20.38 224.82 13.62% 22.90 229.86 15.51% 23.63 135.09 15.56% 22.12 207.68 14.87%
RPMixer 3.2M 16.49 26.75 9.75% 18.82 30.56 11.58% 21.18 35.10 13.46% 18.46 30.13 11.34%
CA HL 30.72 46.96 20.43% 51.56 76.48 37.22% 89.31 125.71 76.80% 54.10 78.97 41.61%
LSTM 98K 19.01 31.21 13.57% 26.49 42.54 20.62% 38.41 60.42 31.03% 26.95 43.07 21.18%
DCRNN 373K 17.52 28.18 12.55% 21.72 34.19 16.56% 28.45 44.23 23.57% 21.81 34.35 16.92%
STGCN 4.5M 19.14 32.64 14.23% 21.65 36.94 16.09% 24.86 42.61 19.14% 21.48 36.69 16.16%
GWNET 469K 16.93 27.53 13.14% 21.08 33.52 16.73% 27.37 42.65 22.50% 21.08 33.43 16.86%
STGODE 1.0M 17.59 31.04 13.28% 20.92 36.65 16.23% 25.34 45.10 20.56% 20.72 36.65 16.19%
1NN - 21.88 36.67 15.19% 25.94 43.36 19.19% 31.29 52.52 25.65% 25.76 43.10 19.43%
Linear 3.5K 19.82 32.39 14.73% 25.20 40.97 19.24% 29.80 49.34 23.09% 24.32 39.88 18.52%
TSMixer 9.5M 18.40 106.28 14.30% 19.77 73.98 15.30% 22.56 87.56 17.80% 19.86 90.20 15.79%
RPMixer 7.8M 15.90 26.08 11.69% 17.79 29.37 13.23% 19.93 33.18 15.11% 17.50 28.90 13.03%
5.3 Benchmark Result
The benchmark results are summarized in Table 1. Our discussion on the performance differences among various methods will proceed in three parts: firstly, we will discuss the 11 baselines included in the original LargeST benchmark Liu et al. (2023); next, we will address the three new time series forecasting baselines we added to the benchmark; finally, we will highlight the performance of our proposed RPMixer method in comparison to the baselines.
[…]
For smaller datasets such as SD and GBA, RPMixer surpasses the performance of baseline methods in the later time steps of the predictions. When it comes to larger datasets like GLA and CA, RPMixer outperforms all baseline methods across every performance measure and time horizon. Even though the proposed method has a high number of parameters similar to TSMixer, it is not overfitting to the loss function and performs exceptionally well across the board. The memory complexity of RPMixer scales linearly with respect to the number of nodes, enabling it to be applied to the largest dataset (CA).
5.4 Ablation Study
We conducted an ablation study on the three most crucial design decisions associated with our proposed RPMixer method. The three decisions were: 1) implementing the identity mapping connection with pre-activation, 2) introducing random projection layers, and 3) processing the time series in the frequency domain. The results of the ablation study are presented in Fig. 7.
Figure 7: The ablation study result measured using average MAE, RMSE, and MAPE. Removing each component resulted in a degradation of the performance.
The pre-activation design emerged as the most consequential design decision. Upon its replacement with a post-activation design, we effectively removed the most important element of the original design, leading to a significant performance decrease. This decline is likely due to the elimination of identity mapping connections from the network. These connections are vital for the ensemble interpretation of the residual connection. Without them, the inclusion of a random projection layer no longer makes sense, as it weakens the ensemble-like-like behavior of the residual blocks (i.e., the mixer blocks in RPMixer), which was the motivation behind the design of the random projection layer.
The random projection layer design was the second most important decision in terms of performance difference. We disabled the random projection layer by converting it into a regular linear layer with trainable weights. This layer helps each mixer block in focusing on different aspects of the inter-node relationship, so its importance is not surprising. To verify our claim that the random projection layer promotes diversity, we constructed a correlation-error diagram comparing the proposed method with the variant without random projection layer. This diagram serves a similar purpose to the kappa-error diagram Margineantu and Dietterich (1997) used for analyzing base learners in an ensemble model. Before discussing the diagram, let us first introduce what a correlation-error diagram is.
Figure 8: The correlation-error diagram illustrates the interplay of diversity and error (MAE) across individual blocks in the model.
A correlation-error diagram is a visualization tool akin to the kappa-error diagram Margineantu and Dietterich (1997), used for analyzing the trade-off between performance and diversity in ensemble models. Each dot in the diagram represents a pair of base learners, with the $y$-axis showing the average performance of the pair’s outputs and the $x$-axis indicating the degree of agreement between them. Specifically, we use the Pearson correlation coefficient to measure agreement between the outputs of a pair of mixer blocks (i.e., the base learners). We use MAE, RMSE, and MAPE as performance measures. The figure created using MAE and Pearson correlation coefficient is shown in Fig. 8. Figures using other performance measures are included in Appendix A.5. We observe that the dots associated with the proposed method with a random projection layer have a wider distribution (larger range in $y$-axis) compared to the variant without the random projection layer. This confirms that our random projection layer effectively increases the diversity of the intermediate representation.
The decision to process the time series in the frequency domain had the least impact on performance. The performance improvement is minor compared to the other two design choices. One possible reason is that the Fourier transformation is a linear transformation, so the majority of its benefits could be learned by the model during training. We also analyze the effect of different design choices at various time horizons and the findings are reported in Appendix A.5.
5.5 Parameter Sensitivity Analysis
In our parameter sensitivity analysis, we focus on two hyper-parameters: the number of mixer blocks and the number of neurons in the random projection layer. We report performance on both the validation and test sets to demonstrate the generalizability of our findings across different data partitions. Fig. 8(a) illustrates the parameter sensitivity analysis for the number of mixer blocks.
(a) Setting the number of block to eight generally yields better results. This observation holds true for both validation and test data.
(b) The number is set by $m_{neuron}\sqrt{n}$, where $n$ is the number of nodes in the graph and $m_{neuron}$ is a factor. Setting this factor to one often has the best performance. This observation holds true for both validation and test data.
Figure 9: The parameter sensitivity analysis for the number of block and number of neurons parameter measured using average MAE, RMSE, and MAPE.
Setting the number of blocks to eight generally yields the best performance on both the validation and test data. Reducing the number of blocks to two significantly diminishes performance, indicating the necessity of multiple blocks to achieve ensemble-like behavior. Increasing the number of blocks to 16 results in minor performance improvement on some datasets. However, as it doubles the model size, the minor gain in performance may not justify the increased computational cost. Similar trends are observed on both the validation and test data, suggesting that we can use the validation set to tune this hyper-parameter.
For the number of neurons in the random projection layer, we set it as a function with respect to the number of nodes in the spatial-temporal data, as more neurons may be required for datasets with larger number of nodes. If the graph has $n$ nodes, we set the number of neurons in the random projection layer as $m_{neuron}\sqrt{n}$, where $m_{neuron}$ is the hyper-parameter controlling the number of neurons in the layer. This design allows the model to automatically use more neurons for datasets with more number of nodes, even if the hyper-parameter $m_{neuron}$ is set to the same value for all datasets. Fig. 8(b) displays the results of the parameter sensitivity analysis for the hyper-parameter $m_{neuron}$.
We evaluated the model under three different settings of $m_{neuron}$: $\{1.0,1.5,2.0\}$. Our observation indicates that setting the hyper-parameter to 1.0 generally leads to good performance, although for some datasets, a setting of 2.0 may yield better results. However, the performance difference is minimal and may not justify the additional computational cost associated with the 2.0 setting. Additionally, the trends on the validation data closely mirror those on the testing data, justifying the use of validation data for setting the hyper-parameter.
6 Conclusion
In this paper, we proposed RPMixer, an all-MLP mixer model that incorporates random projection layers. These layers enhance the diversity among the outputs of each mixer block, thereby improving the overall performance of the model. Our experiments demonstrated that the random projection layers not only improve the diversity of the intermediate representation but also boost the model’s overall performance on large-scale spatial-temporal benchmark datasets in the literature Liu et al. (2023). This work contributes a new approach to spatial-temporal forecasting, potentially providing more accurate and reliable predictions for real-world applications such as traffic flow prediction. In future work, we plan to investigate the potential of applying time series foundation models Yeh et al. (2023a) for tackling spatial-temporal forecasting problems.
Appendix A Supplementary
In this section, we provide supplementary materials for the paper. These materials encompass additional details and results that were excluded from the primary text due to space constraints.
A.1 Complex Linear Layer
As depicted in Fig. 10, the time series in the large spatial-temporal forecasting benchmark dataset exhibits periodicity.
Figure 10: The time series for three randomly selected nodes from the LargeST dataset spans a period of four weeks.
Figure 11: The complex linear layer.
To leverage this periodicity, we opt to process the time series in the frequency domain using a complex linear layer Trabelsi et al. (2018). The computational graph for the complex linear layer is illustrated in Fig. 11. Initially, the input data is converted to the frequency domain using the Fast Fourier Transform (FFT) method. Subsequently, the real and imaginary parts are processed by two distinct linear layers, one containing the real part of the model weights and the other containing the imaginary part. The outputs from these two linear layers are then combined, and finally an inverse FFT layer is used to convert back to the time domain.
The design rationale stems from the fact that, unlike regular linear layers, the weights for the complex linear layer are complex numbers. Consider a simplified scenario where we aim to multiply input data $x$ with a complex weight matrix $W_{\text{real}}+{i\mkern 1.0mu}W_{\text{imag}}$. As depicted in Fig. 11, the input data $x$ is first converted to the frequency domain as $x_{\text{real}}+{i\mkern 1.0mu}x_{\text{imag}}$ using the FFT method. Next, we multiply $x_{\text{real}}+{i\mkern 1.0mu}x_{\text{imag}}$ with $W_{\text{real}}+{i\mkern 1.0mu}W_{\text{imag}}$. Simple algebraic manipulation reveals that the result of this multiplication is
$(W_{\text{real}}x_{\text{real}}-W_{\text{imag}}x_{\text{imag}})+{i\mkern 1.0mu}(W_{\text{real}}x_{\text{imag}}+W_{\text{imag}}x_{\text{real}})$ (6)
This operation is captured in the design illustrated in Fig. 11.
A.2 Dataset Statistics
The statistics for the SD, GBA, GLA, and CA dataset are summarized in Table 2.
Table 2: Dataset statistics.
Data # of nodes # of time steps time range
SD 716 35,040 [1/1/2019, 1/1/2020)
GBA 2,352 35,040 [1/1/2019, 1/1/2020)
GLA 3,834 35,040 [1/1/2019, 1/1/2020)
CA 8,600 35,040 [1/1/2019, 1/1/2020)
A.3 Baseline Method
The details about the 14 baseline methods are provided below:
  • HL: The prediction for all future time steps is generated by using the last value from the historical data.
  • LSTM Hochreiter and Schmidhuber (1997): The Long Short-Term Memory (LSTM), a variant of the recurrent neural network architecture, is specifically designed to process sequential data, making it a general method for time series forecasting. When deploying this architecture on the LargeST dataset, the same model weights are utilized across the time series of different nodes.
  • ASTGCN Guo et al. (2019): The Attention-based Spatial-Temporal Graph Convolutional Network (ASTGCN) model employs a spatial-temporal attention mechanism to capture spatial-temporal correlations. It also uses graph convolutions and conventional convolutions to extract spatial and temporal patterns, respectively.
  • DCRNN Li et al. (2017): The Diffusion Convolutional Recurrent Neural Network (DCRNN) incorporates diffusion convolution, the sequence-to-sequence architecture, and the scheduled sampling technique to capture spatial-temporal patterns.
  • AGCRN Bai et al. (2020): The Adaptive Graph Convolutional Recurrent Network (AGCRN) extends the design of recurrent networks with a node adaptive parameter learning module, enabling the capture of node-specific patterns. Additionally, it includes a data adaptive graph generation module to infer the interdependencies among different nodes.
  • STGCN Yu et al. (2017): The Spatial-Temporal Graph Convolutional Networks (STGCN) employs both graph and temporal convolutions to model the spatial-temporal correlations within the data.
  • GWNET Wu et al. (2019): The Graph WaveNet (GWNET) is another network that relies on convolution operations. Specifically, it employs a gated temporal convolution module and a graph convolution layer to model the spatial-temporal correlation.
  • STGODE Fang et al. (2021): The Spatial-Temporal Graph Ordinary Differential Equation Networks (STGODE) models spatial-temporal dynamics using tensor-based ordinary differential equations. The model design also incorporates a semantical adjacency matrix and temporal dilated convolution modules.
  • DSTAGNN Lan et al. (2022): The Dynamic Spatial-Temporal Aware Graph Neural Network (DSTAGNN) is a method that does not use a predefined static adjacency matrix. Instead, it learns the dynamic spatial associations among nodes and utilizes a spatial-temporal attention module based on multi-order Chebyshev polynomials to capture these associations. To model the temporal associations, it employs gated convolution modules.
  • DGCRN Li et al. (2023): The Dynamic Graph Convolutional Recurrent Network (DGCRN) model combines graph convolution networks with recurrent networks. In this model, a dynamic adjacency matrix is progressively generated by a hyper-network in synchronization with the recurrent steps. This dynamic adjacency matrix, in conjunction with the predefined static adjacency matrix, is utilized to generate predictions.
[…]
  • 1NN: The 1-Nearest-Neighbor (1NN) method serves as a simple baseline for a range of time series problems Bagnall et al. (2017); Dau et al. (2019). We have implemented this baseline by leveraging the matrix profile for its efficiency Yeh et al. (2016); Zimmerman et al. (2019). Notably, the version of the 1NN method benchmarked in this study is the most basic form, where each node is treated as independent from the others. While more advanced techniques from the literature, such as those presented in Yeh et al. (2017); Martínez et al. (2019); Yeh et al. (2022b), could potentially enhance the method, we plan to explore these techniques in our future work.
  • Linear: Having proven its effectiveness in general time series forecasting Zeng et al. (2023), the linear model has been incorporated into our benchmark experiments. Similar to LSTM, the same model weights are used across the time series of various nodes.
  • TSMixer Chen et al. (2023): The Time Series Mixer (TSMixer) is a stacked Multi-Layer Perceptron (MLP) that efficiently extracts information by utilizing mixing operations across both the time and feature dimensions (i.e., nodes for spatial-temporal data). These mixing operations are capable of capturing the relationships between different time steps and nodes.
A.4 Implementation Detail
The experiments were carried out using Python 3.10.11 on a Linux server equipped with an AMD EPYC 7713 64-Core Processor and NVIDIA Tesla A100 GPU. Our implementation leverages PyTorch 2.0.1 to realize the proposed RPMixer, TSMixer, and the Linear model. All three models are trained using the AdamW optimizer Loshchilov and Hutter (2017) with default hyper-parameter settings. The MAE loss function was used, following Liu et al. (2023). In the case of the proposed method, we configure the number of mixer blocks to eight and set the random projection dimension to $\sqrt{n}$, where $n$ is the number of nodes in the graph. For TSMixer, the number of mixer blocks is also set to eight, and the number of hidden dimensions is fixed at 64, in accordance with the parameterization used by the original author in the Traffic prediction dataset experiment Chen et al. (2023). The 1NN method is implemented using pyscamp 0.4.0 Zimmerman et al. (2019). Further details on our implementation can be found in the released source code The Author(2023) (s).
(a) RMSE
(b) MAPE
Figure 12: The correlation-error diagram illustrates the interplay of diversity and error (RMSE and MAPE) across individual blocks in the model.
A.5 Ablation Study
We have produced correlation-error diagrams employing all three measurements (MAE, RMSE, and MAPE). In Section 5.4, we showcase the correlation-error diagram with MAE in Fig. 8. This section introduces the remaining two correlation-error diagrams, namely, the correlation-error diagram with RMSE in Fig. 11(a), and the correlation-error diagram with MAPE in Fig. 11(b). The conclusions remain consistent. The random projection layer aids the proposed method in achieving greater diversity.
Figure 13: The ablation study result measured using per-step MAE, RMSE, and MAPE. The random projection layer benefits the prediction for later time steps more compared to earlier time steps.
To understand the effect of different design choices at various time horizons, we visualized the performance of different variants of the proposed model over the prediction time steps in Fig. 13. For the pre-activation and frequency domain choices, the positive effect is evenly distributed across different time steps. For the random projection layer, the benefits lean more towards longer-term predictions. A possible explanation is that the variance for longer-term prediction is typically higher. Since the random projection layer helps different mixer blocks focus on different dimensions of the multivariate time series, it aids the model in better capturing the later, harder-to-predict time steps.
A.6 Long-Term Time Series Forecasting
The effectiveness of the proposed method in spatial-temporal forecasting tasks has been demonstrated. However, as a general time series forecasting method, we sought to assess its performance against other forecasting methods. To this end, we evaluated our method on seven multivariate long-term time series forecasting datasets (i.e., ETTh1, ETTh2, ETTm1, ETTm2, Weather, Electricity, and Traffic), and compared it with two state-of-the-art methods, TSMixer Chen et al. (2023) and PatchTST Nie et al. (2022), which have shown superior performance over alternatives such as Autoformer Wu et al. (2021), Informer Zhou et al. (2021), TFT Lim et al. (2021), FEDformer Zhou et al. (2022), and linear models Zeng et al. (2023).
The experimental setup followed the guidelines outlined in Nie et al. (2022); Chen et al. (2023). We evaluated the models under long-term time series forecasting settings, with prediction lengths of 96, 192, 336, and 720. The evaluation metrics employed were mean square error (MSE) and mean absolute error (MAE). The dataset was divided into training, validation, and testing subsets, as suggested in Nie et al. (2022); Chen et al. (2023); Zeng et al. (2023). The input length was set to 512 as per Chen et al. (2023), and other hyper-parameters were determined based on the validation set results. The settings for the hyper-parameters are provided with the source code, which can be downloaded from The Author(2023) (s). We employed the AdamW optimizer Loshchilov and Hutter (2017) with the objective of minimizing the mean square error. The results are presented in Table 3.
Table 3: Performance comparisons. We bold the best-performing results.
Data Method Horizon 96 Horizon 192 Horizon 336 Horizon 720
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 PatchTST 0.370 0.400 0.413 0.429 0.422 0.440 0.447 0.468
TSMixer 0.361 0.392 0.404 0.418 0.420 0.431 0.463 0.472
RPMixer 0.444 0.444 0.488 0.475 0.521 0.498 0.625 0.574
ETTh2 PatchTST 0.274 0.337 0.341 0.382 0.329 0.384 0.379 0.422
TSMixer 0.274 0.341 0.339 0.385 0.361 0.406 0.445 0.470
RPMixer 0.173 0.284 0.210 0.317 0.246 0.346 0.340 0.410
ETTm1 PatchTST 0.293 0.346 0.333 0.370 0.369 0.392 0.416 0.420
TSMixer 0.285 0.339 0.327 0.365 0.356 0.382 0.419 0.414
RPMixer 0.358 0.385 0.397 0.406 0.439 0.433 0.503 0.475
ETTm2 PatchTST 0.166 0.256 0.223 0.296 0.274 0.329 0.362 0.385
TSMixer 0.163 0.252 0.216 0.290 0.268 0.324 0.420 0.422
RPMixer 0.111 0.224 0.139 0.252 0.168 0.276 0.212 0.311
Weather PatchTST 0.149 0.198 0.194 0.241 0.245 0.282 0.314 0.334
TSMixer 0.145 0.198 0.191 0.242 0.242 0.280 0.320 0.336
RPMixer 0.149 0.206 0.198 0.250 0.258 0.295 0.343 0.353
Electricity PatchTST 0.129 0.222 0.147 0.240 0.163 0.259 0.197 0.290
TSMixer 0.131 0.229 0.151 0.246 0.161 0.261 0.197 0.293
RPMixer 0.130 0.229 0.149 0.247 0.166 0.264 0.201 0.302
Traffic PatchTST 0.360 0.249 0.379 0.256 0.392 0.264 0.432 0.286
TSMixer 0.376 0.264 0.397 0.277 0.413 0.290 0.444 0.306
RPMixer 0.394 0.277 0.406 0.282 0.415 0.286 0.451 0.307
The results suggest that our proposed method performs comparably to both TSMixer and PatchTST, as confirmed by a $t$-test with $\alpha=0.05$. This implies that the proposed method also achieves state-of-the-art performance on the long-term time series forecasting tasks. Chen et al. Chen et al. (2023) noted that the cross-variate information might not be as significant in these seven datasets. Therefore, the fact that our proposed method achieves state-of-the-art performance indicates that its components for capturing cross-variate information do not degrade performance, even when the cross-variate information is not relevant for the task. This suggests that our method is robust against irrelevant cross-variate information.